### v2

@@ Appendix A Examples
A.1 Connectivity of equilibria with bounded linear utilities
We first investigate the (dis)connectivity of competitive equilibria with bounded linear utilities. We begin by demonstrating that the thrifty competitive equilibria with bounded linear utilities might not be connected.
Example A.1 (Disconnectivity of thrifty competitive equilibrium).
@@ A.1 Connectivity of equilibria with bounded linear utilities
There are two agents and two goods. Each agent has a unit budget. The first agent’s utility is $u_{1}(x_{1})=1.3\min\{x_{11},0.8\}+0.45x_{12}$ and the second agent’s utility is $u_{2}(x_{2})=0.01\min\{x_{21},0.3\}+2\min\{x_{22},0.8\}$.
First, we show there are at least two equilibria. The first equilibrium is with $\mathbf{p}=(1,1)$. Under this price, $x_{11}=0.8$, $x_{12}=0.2$, $x_{21}=0.2$, $x_{22}=0.8$. Another competitive equilibrium is with $\mathbf{p}=(1.3,0.45)$. Under this price, $x_{11}=0.7$, $x_{12}=0.2$, $x_{21}=0.3$, $x_{22}=0.8$.
@@
Now, we show the second competitive equilibrium is isolated under the thrifty assumption.
@@
First, we show the price is isolated. Suppose there is a nearby price $\mathbf{p}^{\prime}=(1.3+\epsilon_{1},0.45+\epsilon_{2})$. For this price, the thrifty demand of the second agent remains the same: $x_{21}=0.3$, $x_{22}=0.8$. To maintain equilibrium, the first agent should receive the same allocation. As $1.3\times 0.7+0.45\times 0.2=1$, $\epsilon_{1}$ and $\epsilon_{2}$ can not be both positive or negative. However, if one $\epsilon$ is negative, the first agent will prefer this item more at $\mathbf{p}^{\prime}$. This implies the allocation of the other item of the first agent is 00. This leads to a contradiction.
@@
Second, we show the allocation is isolated. Suppose there is a nearby equilibrium allocation of the second agent $x^{\prime}_{2}=(x_{21}-\epsilon_{1},x_{22}-\epsilon_{2})$ with at least one of $\epsilon_{1}$ and $\epsilon_{2}$ is strictly positive. In this case, both agents should spend all their budgets as they are non-satiable. Therefore, the equilibrium prices satisfy $p^{\prime}_{1}+p^{\prime}_{2}=2$. For the first agent, the allocation is $x^{\prime}_{11}=0.7+\epsilon_{1}$, $x^{\prime}_{12}=0.2+\epsilon_{2}$. Note that both $\epsilon_{1}$ and $\epsilon_{2}$ are very tiny. This implies $p^{\prime}_{1}:p^{\prime}_{2}=1.3:0.45$. Therefore, $p^{\prime}_{1}=\frac{2}{1.3+0.45}\times 1.3$ and $p^{\prime}_{2}=\frac{2}{1.3+0.45}\times 0.45$. However, the first agent cannot afford these allocations under these prices.
@@
The next example shows that competitive equilibria are not connected without the thrifty assumption.
Example A.2 (Disconnectivity of competitive equilibria).
@@
There are three agents and three goods. Each agent has a unit budget. The first agent’s utility is $u_{1}(x_{1})=1.3\min\{x_{11},0.8\}+0.45x_{12}$, the second agent’s utility is $u_{2}(x_{2})=0.01\min\{x_{21},0.3\}+2\min\{x_{22},0.8\}+10^{-10}x_{23}$, and the third agent’s utility is $u_{3}(x_{3})=x_{33}$.
@@
First, we show there are at least two equilibrium. The first equilibrium is with $\mathbf{p}=(1,1,1)$, and the allocation is $x_{1}=(0.8,0.2,0)$, $x_{2}=(0.2,0.8,0)$, and $x_{3}=(0,0,1)$. The second equilibrium is with $\mathbf{p}=(1.3,0.45,1.25)$ and the allocation is $x_{1}=(0.7,0.2,0)$, $x_{2}=(0.3,0.8,0.2)$, and $x_{3}=(0,0,0.8)$.
@@
First, we demonstrate that the price is isolated. Suppose there is a nearby equilibrium price $\mathbf{p}^{\prime}=(1.3+\epsilon_{1},0.45+\epsilon_{2},1.25+\epsilon_{3})$. For this price, the allocation of the second agent on the first two items is the same. Therefore, the first agent should receive the same allocation. As $1.3\times 0.7+0.45\times 0.2=1$, $\epsilon_{1}$ and $\epsilon_{2}$ can not be both positive or negative. Similar to the previous example, if one of $\epsilon_{1}$ and $\epsilon_{2}$ is negative, then the first agent will prefer this item more at this nearby price. This implies the allocation of the other item of the first agent is 00. This leads to the conclusion $\epsilon_{1}=\epsilon_{2}=\epsilon_{3}=0$.
Second, we show the allocation is isolated. Suppose there is a nearby equilibrium allocation $\mathbf{x}^{\prime}$ with the equilibrium $\mathbf{p}^{\prime}$. As $x_{11}<0.8$, $p^{\prime}_{1}:p^{\prime}_{2}=1.3:0.45$.  [^2] Otherwise, the allocation of the first agent will be changed significantly. Additionally, $p^{\prime}_{1}\times x^{\prime}_{11}+p^{\prime}_{2}\times x^{\prime}_{12}=1$. This implies $p_{1}^{\prime}$ is very close to $p_{1}$ and $p_{2}^{\prime}$ is very close to $p_{2}$: if $\|x^{\prime}-x\|_{\infty}\leq\epsilon$, then $|p_{1}^{\prime}-p_{1}|\leq O(\epsilon)$ and $|p_{2}^{\prime}-p_{2}|\leq O(\epsilon)$.  [^3] Note that $p^{\prime}_{1}:p^{\prime}_{2}=1.3:0.45=p_{1}:p_{2}$. By budget constraint, $\frac{1.3}{0.45}p^{\prime}_{2}x^{\prime}_{11}+p^{\prime}_{2}x^{\prime}_{12}=\frac{1.3}{0.45}p_{2}x_{11}+p_{2}x_{12}$. Therefore, $p_{2}^{\prime}-p_{2}=\frac{p^{\prime}_{2}(\frac{1.3}{0.45}(x_{11}-x^{\prime}_{11})+(x_{12}-x^{\prime}_{12}))}{\frac{1.3}{0.45}x_{11}+x_{12}}$. We have a similar result for $p_{1}^{\prime}-p_{1}$. This also implies $|p_{3}^{\prime}-p_{3}|\leq O(\epsilon)$ as $p^{\prime}_{1}+p^{\prime}_{2}+p^{\prime}_{3}=3$. Since price is isolated, $p^{\prime}_{1}=p_{1}$, $p^{\prime}_{2}=p_{2}$, and $p^{\prime}_{3}=p_{3}$. Under this price, $\mathbf{x}$ is the only allocation.
A.2 Non-monotonicity of Gale Demands
@@
Example A.3 (Non-monotonicity of Gale demands).
There are two goods and a single agent $i$ with $b_{i}=1$. The agent’s utility function is $u_{i}(x_{i})=\min\{x_{i1}+x_{i2},2+0.1\cdot x_{i1}+0.2\cdot x_{i2}\}$. When $\mathbf{q}=\left(\frac{0.1}{2/0.9+0.2\cdot 0.01},\frac{0.2}{2/0.9+0.2\cdot 0.01}\right)$, the associated Gale demand is $\mathcal{GD}_{i}(b_{i},\mathbf{q})=\{(2/0.9,0.01)\}$ and the utility is $u_{i}(\mathcal{GD}_{i}(b_{i},\mathbf{q}))=2/0.9+0.2\cdot 0.01$. When $\mathbf{q}^{\prime}=\left(\frac{1}{0.01+2/0.85},\frac{0.2}{0.01+2/0.85}\right)$, the Gale demand becomes $\mathcal{GD}_{i}(b_{i},\mathbf{q}^{\prime})=\{(0.01,2/0.85)\}$ and $u_{i}(\mathcal{GD}_{i}(b_{i},\mathbf{q}^{\prime}))=2/0.85+0.01$. Clearly, $\mathbf{q}<\mathbf{q}^{\prime}$, yet $u_{i}(\mathcal{GD}_{i}(b_{i},\mathbf{q}))<u_{i}(\mathcal{GD}_{i}(b_{i},\mathbf{q}^{\prime}))$.
A.3 Tight Example on Price of Anarchy
The next example shows that the bound in Theorem 1.1 is tight.
Example A.4 (Tight Example on Price of Anarchy).
We consider the case with two agents and two goods. The first agent has a budget of $1$, and desires the first good: $u_{1}(x_{1})=x_{11}$. The second agent has a budget of $\mathrm{e}-1$ and has a utility function $u_{2}(x_{2})=\varepsilon x_{21}+\min\{x_{22},1\}$, where $\varepsilon>0$ is sufficiently small. There is a competitive equilibrium with prices $p_{1}=\mathrm{e}$ and $p_{2}=0$, and allocations $x_{1}=(\frac{1}{\mathrm{e}},0)$ and $x_{2}=(\frac{\mathrm{e}-1}{\mathrm{e}},1)$. The Nash welfare objective for this allocation is $\left(\frac{1}{\mathrm{e}}(1+\varepsilon\cdot\frac{\mathrm{e}-1}{\mathrm{e}})^{\mathrm{e}-1}\right)^{\frac{1}{\mathrm{e}}}$. In the allocation maximizing Nash welfare, we have $y_{1}=(1,0)$ and $y_{2}=(0,1)$, where the Nash welfare objective is $1$. When $\varepsilon\rightarrow 0$, the difference of these two scenarios converges to $\left(\frac{1}{\mathrm{e}}\right)^{\frac{1}{\mathrm{e}}}$.
A.4 Utility Difference in Different Equilibria
The following example shows agents might receive significantly different utilities in different equilibria.
Example A.5 (Utility Difference in Different Equilibria).
Consider a scenario involving $n$ agents and $n$ goods, where each agent has a unit budget. For $11\leq i\leq n-$ 1, agent $i$’s utility is defined as $u_{i}(x_{i})=\min\{x_{ii},1\}+\epsilon x_{in}$, while agent $n$’s utility is given by $u_{i}(x_{i})=x_{nn}$. There exist at least two competitive equilibria. In the first competitive equilibrium, all goods are priced at $1$, and each good $i$ is allocated to agent $i$ for $1\leq i\leq n$, resulting in a utility of $1$ for agent $n$. However, in the second competitive equilibrium, prices are set to 00 for the first $n-1$ goods and the price of the $n$-th good is $n$. Consequently, goods $1$ to $n-1$ are allocated to agents $1$ to $n-1$ respectively, while the $n$-th good is uniformly allocated to all agents. In this equilibrium, agent $n$ can only attain a utility of $\approx\frac{1}{n}$.
A.5 Tight Example on Demand-Approximate CE
The following example shows Theorem 1.2 is tight.
Example A.6 (Tight Example on Demand-Approximate CE).
Consider the case with two agents and one good. Both agents have unit budgets. The first agent’s utility on this good is linear: $u_{1}(x_{1})=x_{1}$. The second agent’s utility is defined as follows: $u_{2}(x_{2})=\alpha x_{2}$ for $x_{2}\leq\frac{1}{\alpha}$ and $u_{2}(x_{2})=1+\beta(x_{2}-\frac{1}{\alpha})$ for $x_{2}>\frac{1}{\alpha}$, where $\alpha$ is a large number and $\beta\leq\alpha$. When $\beta=\frac{1-\frac{1}{\alpha}}{1-\frac{2}{\alpha}}$, the Nash welfare maximizing allocation is $x_{1}=1-\frac{1}{\alpha}$ and $x_{2}=\frac{1}{\alpha}$, with the dual variable $q$ equal to $\frac{1}{x_{1}}=\frac{1}{1-\frac{1}{\alpha}}$. Under this price, agent $2$ can obtain $x^{\prime}_{2}=1-\frac{1}{\alpha}$ with his/her budget, resulting in a utility of $1+\beta(x^{\prime}_{2}-\frac{1}{\alpha})=2-\frac{1}{\alpha}$. In contrast, in the Nash welfare maximizing allocation, agent $2$ only receives a utility of $1$.
A.6 (In)-approximate Nash Welfare utility by Equilibrium utility
@@ A.5 Tight Example on Demand-Approximate CE
In this paper, we show for $\Sigma$-Gale-substitues utilities, the utility of any agent at the Nash welfare maximizing allocation is at least half the maximum utility they may get at any CE. It’s natural to ask, inversely, if the utility agent can attain at any CE can approximate the utility they receive at Nash welfare maximizing allocation. We show a constant approximation is not achievable by providing an example in which the CE is unique and there exists an agent whose utility at this CE is much worse than his/her utility at Nash welfare maximizing allocation.
Example A.7.
Consider a scenario with $n+2$ agents and $n+1$ goods. Set $\epsilon<1$. For $1\leq i\leq n$, let $u_{i}(x_{i})=\min\{x_{ii},0.5\}+\epsilon x_{i(n+1)}$. Additionally, set $u_{n+1}(x_{n+1})=x_{(n+1)(n+1)}$ and $u_{n+2}(x_{n+2})=\sum_{j=1}^{n}x_{ij}$. We first demonstrate the unique competitive equilibrium is achieved when $p_{j}=\frac{2}{n}$ for $1\leq j\leq n$ and $p_{n+1}=n$. In this equilibrium, $x_{ii}=0.5$ and $x_{i(n+1)}=\frac{n-1}{n^{2}}$ for $1\leq i\leq n$; $x_{(n+1)(n+1)}=\frac{1}{n}$; and $x_{(n+2)j}=0.5$ for $1\leq j\leq n$.
@@
We now show this is the only competitive equilibrium. First, for all $j:1\leq j\leq n$, $p_{j}>0$. Otherwise, agent $n+2$ would receive infinite goods. Given $p_{j}>0$ for any $1\leq j\leq n$, $x_{(n+2)j}\geq 0.5$. Consequently, the first $n$ goods share the same prices and $10.5p_{j}n\leq\sum_{j}x_{(n+2)j}p_{j}=$ 1 which implies $p_{j}\leq\frac{2}{n}$. If $p_{j}<\frac{2}{n}$ for $1\leq j\leq n$, then $p_{n+1}=\sum_{i}b_{i}-\sum_{j}p_{j}>n$. In this case, $x_{ii}=0.5$ for $1\leq i\leq n$ and $10.5p_{j}n=$ 1 for $1\leq j\leq n$, which implies the result.
Therefore, agent $n+1$ receives utility $\frac{1}{n}$ in the unique competitive equilibrium. However, for a small enough $\epsilon$, this agent receives $1$ in the maximal Nash welfare solution.
A.7 (In)-approximation for general utilities
@@
Example A.8.
There are three types of agents (A, B, C): $1$ type-A agent; $n^{2}$ type-B agents; $n$ type-C agents; and four types of goods (O, P, R, S): $1$ type-O good; $n$ type-P goods; $n^{2}$ type-R goods; $n$ type-S goods. In total, there are $N=n^{2}+n+1$ agents and $M=n^{2}+2n+1$ goods.
@@
The type-A agent is interested in the type-O good and type-P goods. Type-A agent has a bounded linear utility with equal preferences over these type-O and type-P goods: $v_{ij}=1$ for $i=A$ and $j\in O\cup P$. This agent has an upper bound of 1 on the utility of each good: $u_{A}(x_{A})=\sum_{j\in O\cup P}\min\{x_{Aj},1\}$.
For type-B agents, let’s label each of these agents by $(s,t)$ for $s\in[1,n]$ and $t\in[1,n]$ and also label each of type-R goods by $(s,t)$ for $s\in[1,n]$ and $t\in[1,n]$. Agent $(s,t)$ is interested in the corresponding $(s,t)$ good in R and the $s$-th good in P. The utility function of agent $i=(s,t)$ is $\min\{x_{iP_{s}}+\frac{n}{n+1},x_{iR_{s,t}}\}$, where $x_{iP_{s}}$ is agent $i$’s allocation of $P_{s}$ and $x_{iR_{s,t}}$ is agent $i$’s allocation of $R_{s,t}$.
@@
For type-C agents, the $s$-th type-C agent is only interested in the $s$-th type-S good and all type-R good with label $(s,\cdot)$. This agent also has a bounded linear utility with equal preference ($v_{ij}=1$) over these goods and an upper bound, $l_{ij}=1$, on the utility of each good.
We first consider the maximal Nash welfare solution. Here, $k$ is a parameter, which will be specified later.
The type-A agent receives type-O good and $k+\frac{1}{n+1}$ portion of type-P goods, his/her utility will be $1+nk+\frac{n}{n+1}$.
@@
Each type-B agent obtains $1-\frac{k}{n}$ portion of corresponding type-R good and $\frac{1}{n+1}-\frac{k}{n}$ portion of corresponding type-P good, yielding a utility of $1-\frac{k}{n}$.
The type-C agents receive the remaining goods, resulting in a utility of $1+k$.
@@
The goal is to maximize $(1+nk+\frac{n}{n+1})(1-\frac{k}{n})^{n^{2}}(1+k)^{n}$. By calculation, the optimal $k$ is found to be no larger than $n^{-\frac{1}{2}}$.  [^4] First, we take the log and calculate the gradient, which is $\frac{n}{1+nk+\frac{n}{n+1}}+\frac{n}{1+k}-\frac{n}{1-\frac{k}{n}}$. Note that this is a decreasing function with $k$. So, in order to show the optimal $k$ is no larger than $n^{-\frac{1}{2}}$, we only need to show when $k=n^{-\frac{1}{2}}$, $\frac{n}{1+nk+\frac{n}{n+1}}+\frac{n}{1+k}\leq\frac{n}{1-\frac{k}{n}}$. This is true as $\frac{n}{1-\frac{k}{n}}-\frac{n}{1+k}=\frac{(n+1)k}{1+k-\frac{k}{n}-\frac{k^{2}}{n}}\geq nk=\frac{1}{k}\geq\frac{n}{1+nk+\frac{n}{n+1}}$. Therefore, the type-A agent receives at most $1+\sqrt{n}+\frac{n}{n+1}$ utility in the maximal Nash welfare solution.
On the other hand, let’s consider the (thrifty) competitive equilibrium where all agents have unit budgets.

### v3

@@ Appendix E Examples
E.1 Connectivity of equilibria with bounded linear utilities
We first investigate the (dis)connectivity of competitive equilibria with bounded linear utilities. Our initial example illustrates that the set of thrifty competitive equilibria may be disconnected, even when agents have bounded linear utility functions, where “thrifty” denotes that, given the prices, each agent selects the least costly bundle from the set of utility-maximizing bundles.
Example E.1 (Disconnectivity of thrifty competitive equilibrium).
@@ E.1 Connectivity of equilibria with bounded linear utilities
There are two agents and two goods. Each agent has a unit budget. The agents’ utility functions are
[…]
First, we show that there are at least two equilibria. The first equilibrium is
$p_{1}=1\,,\quad p_{2}=1\,,\quad x_{11}=0.8\,,\quad x_{12}=0.2\,,\quad x_{21}=0.2\,,\quad x_{22}=0.8\,.$
Another competitive equilibrium is
$p_{1}=1.3\,,\quad p_{2}=0.45\,,\quad x_{11}=0.7\,,\quad x_{12}=0.2\,,\quad x_{21}=0.3\,,\quad x_{22}=0.8\,.$
@@
We show the second competitive equilibrium is isolated under the thrifty assumption.
@@
First, we show the price is isolated. Suppose there is a nearby equilibrium price $\mathbf{p}^{\prime}=(1.3+\varepsilon_{1},0.45+\varepsilon_{2})$. For this price, the thrifty demand of the second agent remains the same: $x_{21}=0.3$, $x_{22}=0.8$. To maintain equilibrium, the first agent should receive the same allocation. As $1.3\times 0.7+0.45\times 0.2=1$, $\varepsilon_{1}$ and $\varepsilon_{2}$ cannot be both positive or negative. However, if $\varepsilon_{j}<0$, the first agent will prefer good $j$ more at $\mathbf{p}^{\prime}$. This implies the allocation of the other good of the first agent is 0. This leads to a contradiction.
@@
Second, we show the allocation is isolated. Note that the second agent is currently satiated. Suppose there is a nearby equilibrium allocation. Agent 2 must have a different allocation, which can only be smaller. Thus, $\mathbf{x}^{\prime}_{2}=(x_{21}-\varepsilon_{1},x_{22}-\varepsilon_{2})$ with $\varepsilon_{1},\varepsilon_{2}\geq 0$, and at least one of them is strictly positive. In the allocation $\mathbf{x}^{\prime}$, both agents are non-satiated, and therefore they are spending their full budgets. Therefore, the equilibrium prices satisfy $p^{\prime}_{1}+p^{\prime}_{2}=2$. For the first agent, the allocation is $x^{\prime}_{11}=0.7+\varepsilon_{1}$, $x^{\prime}_{12}=0.2+\varepsilon_{2}$. Note that both $\varepsilon_{1}$ and $\varepsilon_{2}$ are very tiny. This implies $p^{\prime}_{1}:p^{\prime}_{2}\approx 1.3:0.45$. Therefore, $p^{\prime}_{1}\approx\frac{2}{1.3+0.45}\times 1.3$ and $p^{\prime}_{2}\approx\frac{2}{1.3+0.45}\times 0.45$. However, the first agent cannot afford these allocations under these prices.
@@
Nevertheless, it remains possible that thrifty competitive equilibria are connected through intermediate equilibria that are not thrifty. However, with a slight modification of the previous example, the next example shows that competitive equilibria are not connected without the thrifty assumption.
Example E.2 (Disconnectivity of competitive equilibria).
@@
There are three agents and three goods. Each agent has a unit budget. The utility functions are
[…]
$u_{3}(\mathbf{x}_{3})$ $=x_{33}\,.$
@@
First, we show there are at least two equilibria. The first equilibrium is with $\mathbf{p}=(1,1,1)$, and the allocation is $\mathbf{x}_{1}=(0.8,0.2,0)$, $\mathbf{x}_{2}=(0.2,0.8,0)$, and $\mathbf{x}_{3}=(0,0,1)$. The second equilibrium is with $\mathbf{p}=(1.3,0.45,1.25)$ and the allocation is $\mathbf{x}_{1}=(0.7,0.2,0)$, $\mathbf{x}_{2}=(0.3,0.8,0.2)$, and $\mathbf{x}_{3}=(0,0,0.8)$.
@@
First, we demonstrate that the price is isolated. Suppose there is a nearby equilibrium price $\mathbf{p}^{\prime}=(1.3+\varepsilon_{1},0.45+\varepsilon_{2},1.25+\varepsilon_{3})$. For this price, the allocation of the second agent on the first two goods is the same. Therefore, the first agent should receive the same allocation. As $1.3\times 0.7+0.45\times 0.2=1$, $\varepsilon_{1}$ and $\varepsilon_{2}$ cannot be both positive or negative. Similar to the previous example, if one of $\varepsilon_{1}$ and $\varepsilon_{2}$ is negative, then the first agent will prefer this good more at this nearby price. This implies the allocation of the other good of the first agent is 0. This leads to the conclusion $\varepsilon_{1}=\varepsilon_{2}=\varepsilon_{3}=0$.
Second, we show the allocation is isolated. Suppose there is a nearby equilibrium allocation $\mathbf{x}^{\prime}$ with the equilibrium $\mathbf{p}^{\prime}$. As $x_{11}<0.8$, $p^{\prime}_{1}:p^{\prime}_{2}=1.3:0.45$. [^7] Otherwise, the allocation of the first agent will be changed significantly. Additionally, $p^{\prime}_{1}\times x^{\prime}_{11}+p^{\prime}_{2}\times x^{\prime}_{12}=1$. This implies $p_{1}^{\prime}$ is very close to $p_{1}$ and $p_{2}^{\prime}$ is very close to $p_{2}$: if $\|\mathbf{x}^{\prime}-\mathbf{x}\|_{\infty}\leq\varepsilon$, then $|p_{1}^{\prime}-p_{1}|\leq O(\varepsilon)$ and $|p_{2}^{\prime}-p_{2}|\leq O(\varepsilon)$. [^8] Note that $p^{\prime}_{1}:p^{\prime}_{2}=1.3:0.45=p_{1}:p_{2}$. By budget constraint, $\frac{1.3}{0.45}p^{\prime}_{2}x^{\prime}_{11}+p^{\prime}_{2}x^{\prime}_{12}=\frac{1.3}{0.45}p_{2}x_{11}+p_{2}x_{12}$. Therefore, $p_{2}^{\prime}-p_{2}=\frac{p^{\prime}_{2}(\frac{1.3}{0.45}(x_{11}-x^{\prime}_{11})+(x_{12}-x^{\prime}_{12}))}{\frac{1.3}{0.45}x_{11}+x_{12}}$. We have a similar result for $p_{1}^{\prime}-p_{1}$. This also implies $|p_{3}^{\prime}-p_{3}|\leq O(\varepsilon)$ as $p^{\prime}_{1}+p^{\prime}_{2}+p^{\prime}_{3}=3$. Since price is isolated, $p^{\prime}_{1}=p_{1}$, $p^{\prime}_{2}=p_{2}$, and $p^{\prime}_{3}=p_{3}$. Under this price, $\mathbf{x}$ is the only allocation.
E.2 Non-monotonicity of Gale Demands
@@
Example E.3 (Non-monotonicity of Gale demands).
Consider a market with two goods and a single agent with a unit budget, $b=1$. The agent’s utility function is
$u(\mathbf{x})=\min\{x_{1}+x_{2},2+0.1\cdot x_{1}+0.2\cdot x_{2}\}.$
Consider the price vector:
$\mathbf{q}=\left(\frac{0.1}{2/0.9+0.2\cdot 0.01},\frac{0.2}{2/0.9+0.2\cdot 0.01}\right).$
At this price, the agent’s Gale demand and corresponding utility are
$\mathcal{GD}(\mathbf{q},b)=\{(2/0.9,0.01)\},\quad u(\mathcal{GD}(\mathbf{q},b))=2/0.9+0.2\cdot 0.01.$
Now, consider a higher price vector:
$\mathbf{q}^{\prime}=\left(\frac{1}{0.01+2/0.85},\frac{1}{0.01+2/0.85}\right).$
Under price $\mathbf{q}^{\prime}$, the agent’s Gale demand and utility become
$\mathcal{GD}(\mathbf{q}^{\prime},b)=\{(0.01,2/0.85)\}\quad u(\mathcal{GD}(\mathbf{q}^{\prime},b))=2/0.85+0.01.$
Clearly, $\mathbf{q}<\mathbf{q}^{\prime}$ component-wise, yet we observe that $u(\mathcal{GD}(\mathbf{q},b))<u(\mathcal{GD}(\mathbf{q}^{\prime},b))$.
It is also worth noting that such a case cannot arise under separable utilities (see the proof of Lemma C.2 for further details).
E.3 (In)-approximate Nash Welfare utility by Equilibrium utility
@@
Theorem 1.3 shows that for $\Sigma$-Gale-substitutes utilities, $u_{i}(\mathbf{y}_{i})\geq\frac{1}{2}u_{i}(\mathbf{x}_{i})$ for the Nash welfare maximizing allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$ and CE allocation $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$. One may ask the converse: namely, for any agent $i$, is there a CE $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$ such that $u_{i}(\mathbf{x}_{i})\geq C\cdot u_{i}(\mathbf{y}_{i})$ for some constant C>0C>0999We note that the utility achieved by an agent in a Nash welfare maximizing allocation is unique, since the logarithm of the Nash welfare is a strictly concave function.? We show a constant approximation is not achievable by providing an example in which the CE is unique and there exists an agent whose utility at this CE is much worse than his/her utility at Nash welfare maximizing allocation.
Example E.4.
Consider a market with $n+2$ agents and $n+1$ goods. Let $\varepsilon<1$ be a small constant. For each agent $1\leq i\leq n$, the utility function is given by:
$u_{i}(\mathbf{x}_{i})=\min\{x_{ii},0.5\}+\varepsilon x_{i(n+1)}.$
For the remaining two agents, define their utilities as:
$u_{n+1}(\mathbf{x}_{n+1})=x_{(n+1)(n+1)},\quad\text{and}\quad u_{n+2}(\mathbf{x}_{n+2})=\sum_{j=1}^{n}x_{(n+2)j}\,.$
We claim that the unique competitive equilibrium occurs with prices:
$p_{j}=\frac{2}{n}\quad\text{for }1\leq j\leq n,\qquad p_{n+1}=n.$
At this equilibrium, the allocation is as follows:
  • for agent $i$ such that $1\leq i\leq n$: $x_{ii}=0.5$ and $x_{i(n+1)}=\frac{n-1}{n^{2}}$;
  • for agent $n+1$: $x_{(n+1)(n+1)}=\frac{1}{n}$;
  • for agent $n+2$: $x_{(n+2)j}=0.5$ for $1\leq j\leq n$.
We now show that this is the only competitive equilibrium.
@@
First, for all $1\leq j\leq n$, $p_{j}>0$, since otherwise agent $n+2$ could obtain an unbounded amount of good $j$, violating feasibility. Given $p_{j}>0$, and since $x_{(n+2)j}\geq 0.5$ for all $1\leq j\leq n$, the first $n$ goods share the same prices, and, additionally, the total spending of agent $n+2$ satisfies: $1=\sum_{j=1}^{n}x_{(n+2)j}p_{j}\geq 0.5\cdot p_{j}\cdot n$, which implies $p_{j}\leq\frac{2}{n}$. We now show $p_{j}=\frac{2}{n}$. If $p_{j}<\frac{2}{n}$, then, given $\sum_{j}p_{j}=\sum_{i}b_{i}$, $p_{n+1}=\sum_{i}b_{i}-\sum_{j=1}^{n}p_{j}>n$. In this case, agents $1\leq i\leq n$ receive $x_{ii}=0.5$, and therefore $x_{(n+2)i}=1-0.5=0.5$ for all $1\leq i\leq n$. This implies $0.5\cdot p_{j}\cdot n=1$, which leads back to $p_{j}=\frac{2}{n}$.
Therefore, this pricing and allocation define the unique competitive equilibrium.
Finally, observe that agent $n+1$ receives utility $\frac{1}{n}$ in this equilibrium. In contrast, under the allocation that maximizes Nash welfare, for sufficiently small $\varepsilon$, agent $n+1$ would receive utility $1$. This illustrates that the utility an individual agent receives in a Nash welfare maximizing allocation can be substantially higher than in the corresponding market equilibrium.
E.4 (In)-approximation for general utilities
@@
Example E.5.
We consider an economy with three types of agents (A, B, C):
11 type-A agent, $n^{2}$ type-B agents, $n$ type-C agents;
and four types of goods (O, P, R, S):
11 type-O good, $n$ type-P goods, $n^{2}$ type-R goods, $n$ type-S goods.
In total, there are $N=n^{2}+n+1$ agents and $M=n^{2}+2n+1$ goods.
@@
The type-A agent is interested in the type-O good and type-P goods with equal preferences: $v_{ij}=1$ for $i=A$ and $j\in O\cup P$. However, the utility from each good is capped at $1$, so the utility is given by:
$u_{A}(\mathbf{x}_{A})=\sum_{j\in O\cup P}\min\{x_{Aj},1\}.$
The type-B agents are indexed by pairs $(s,t)$ where $s,t\in[n]$, for a total of $n^{2}$ agents. Similarly, the type-R goods are indexed by $(s,t)$. Each type-B agent $(s,t)$ is interested in
  • the $s$-th type-P good, and
  • the type-R good with index $(s,t)$.
The utility function of type-B agent with index $i=(s,t)$ is
$u_{i}(\mathbf{x}_{i})=\min\{x_{iP_{s}}+\frac{n}{n+1},x_{iR_{s,t}}\},$
where $x_{iP_{s}}$ is agent $i$’s allocation of the $s$-th type-P good and $x_{iR_{s,t}}$ is agent $i$’s allocation of the type-R good with index $(s,t)$.
@@
The type-C agents are indexed by $s\in[n]$. Each type-C agent with index $s$ is only interested in the $s$-th type-S good and all type-R goods with index $(s,\cdot)$, and the utility from each good is capped at $1$. The utility function type-C agent with index $i=s$ is
$u_{i}(\mathbf{x}_{i})=\sum_{j\in{S_{s}}\cup{R_{(s,\cdot)}}}\min\{x_{ij},1\},$
where $S_{s}$ represents the $s$-th type-S good and $R_{(s,\cdot)}$ represents the set of type-R goods with index $(s,\cdot)$.
The maximal Nash Welfare solution
Let $k$ be a parameter, which will be specified later.
@@
The type-A agent receives the type-O good and $k+\frac{1}{n+1}$ portion of type-P goods. This results in a total utility of $1+nk+\frac{n}{n+1}$. Each type-B agent obtains $1-\frac{k}{n}$ portion of corresponding type-R good and $\frac{1}{n+1}-\frac{k}{n}$ portion of corresponding type-P good, yielding a utility of $1-\frac{k}{n}$. The type-C agents receive the remaining goods, resulting in a utility of $1+k$.
@@
The goal is to maximize $(1+nk+\frac{n}{n+1})(1-\frac{k}{n})^{n^{2}}(1+k)^{n}$. By calculation, the optimal $k$ is at most $n^{-\frac{1}{2}}$. [^10] First, we take the log and calculate the gradient, which is $\frac{n}{1+nk+\frac{n}{n+1}}+\frac{n}{1+k}-\frac{n}{1-\frac{k}{n}}$. Note that this is a decreasing function with $k$. So, in order to show the optimal $k$ is no larger than $n^{-\frac{1}{2}}$, we only need to show when $k=n^{-\frac{1}{2}}$, $\frac{n}{1+nk+\frac{n}{n+1}}+\frac{n}{1+k}\leq\frac{n}{1-\frac{k}{n}}$. This is true as $\frac{n}{1-\frac{k}{n}}-\frac{n}{1+k}=\frac{(n+1)k}{1+k-\frac{k}{n}-\frac{k^{2}}{n}}\geq nk=\frac{1}{k}\geq\frac{n}{1+nk+\frac{n}{n+1}}$. Therefore, the type-A agent receives at most $1+\sqrt{n}+\frac{n}{n+1}$ utility in the maximal Nash welfare solution.
Thrifty competitive equilibrium